OpenStudy (anonymous):

What is the quotient: (3x2 + 8x - 3) ÷ (x + 3)

OpenStudy (anonymous):

3x + 5, r = 1 3x - 5 3x + 1 3x - 1

OpenStudy (anonymous):

\[\frac{3x^2+8x-3}{x + 3}= \frac{(3x-1)(x+3)}{x + 3}\] The (x+3) terms cancel each other out, and you're left with \[3x-1\]